(1 point) A steel ball weighing 128 pounds is suspended from a spring. This stretches the spring 528 feet. The ball is starte(1 point) Find the solution to the linear system of differential equations {s: -17x + 14y -21x + 18y satisfying the initial cPLEASE ANSWER BOTH PROBLEM SETS PLEASE ANSWER BOTH
Answer #1

128 pound and Steel steel ball weight öll m = 128 32 m = 128 uky L=128 feet 577 K = Weight 128 こ 128 577 x=577 X= 4 myl (t) +y (t) = -1 62 (sini2t) tize 4/2 (2 cosizt y(o)=7=1262 is = 7 / 12 y (t) = 7 / 4 th е sin (12t) As. x = -170 +144 y= -21 xele EN] =B] et V₂ = 1, Vi=1 V N eigen vector corresponding ? =4 bem ] [] = [:] -21 -21 14 im -21 14 [*] =[:] nu let m₂ =1, mа e-36 о Filt? .) 11 + беч | 9 € + Чече x (t) = 2e3e y (t) = 20 3t + bent An